# Speculative Diffusion Decoding: Accelerating Language Generation Through Diffusion

### v3

@@ 1 Introduction
As autoregressive language modeling with transformers [1] is scaled to larger compute levels, performance improves and new capabilities emerge [2, 3]. Indeed, scaling has been shown to to improve the performance of large language models (LLMs) for a diverse array of tasks, including code generation, question answering, summarization, and many other use cases [4, 5, 6]. For instance, models such as LLaMA 3.2 90B [6], ChatGPT [7], Cohere 52B [8], Google’s Gemini-ULTRA [9] exemplify the ongoing trend of deploying and releasing increasingly large models, enabling broader access and application across various domains.
However, while these desired capability arise, running LLMs in inference mode for millions of users produces burdensome electricity, time, and monetary demands. Many methods exist to mitigate these costs – including sparsity, quantization, and distillation – but they often introduce new tradeoffs, e.g., their application can degrade the performance of the model [10].
@@
Notably, since both the drafting model’s speed and its alignment relative to the target model are critical to the success of speculative decoding, simultaneous improvements in each of these areas are necessary to ensure speculative decoding’s relevance to future, more capable target models. For instance, a small GPT-2 [13] drafting model could produce drafts that are often rejected by GPT-4 [4], and simply scaling the drafting model to address its weaker generations risks diminishing the speed advantage necessary to speculative decoding’s success. To address this challenge, previous efforts have focused on introducing additional parallelization techniques that incorporate prediction trees and branching to refine the drafting process [14, 15, 16]. However, the gain in generation efficiency are at the expense of much increased number of operations and/or memory for each generation.
This paper proposes a fundamentally different approach to improve speculative decoding: It proposes to replace the auto-regressive drafter with recently introduced discrete diffusion models [17, 18]. These models offer several key advantages when used as drafters: Firstly, they provide a smooth trade off between the compute cost of generation and the quality of generation (via the number of reverse diffusion steps). Second, while they have historically struggled relative to traditional language models, recent diffusion models have been shown to require 32×32$\times$32 × fewer function evaluations than autoregressive models to produce text with comparable perplexity [17], with more recent works reporting even further speedups [18]. This is a trend which is anticipated to continue. Finally, future advances in diffusion model generation quality are highly aligned with their ability to perform strongly as speculative drafters: as drafted tokens are accepted by the target model at a higher rate, a larger number of proposed drafted tokens becomes optimal from an efficiency/speed point of view, and (unlike sequential drafters) diffusion models can easily accommodate generation of many more tokens since they are able to generate entire sequences in a single step.
@@
    The proposed method ensures that all generations from the diffusion language model, which are empirically shown to produce outputs with significantly higher perplexity than current state-of-the-art autoregressive models [18, 17, 19, 20], align with the outputs generated by larger, more computationally demanding models.
@@
    Finally, the paper sets a new benchmark for speed in language completion tasks on the CNN/DM [21]and OpenWebText datasets [22].
@@ 2 Related Work
While autoregressive language models provide state-of-the-art performance on language generation tasks, the incremental decoding used by these architectures results in significant overhead at inference time [23]. This is largely a result of the inability to parallelize the sequential process of generating tokens in the output sequence as each token generation is dependent upon the preceding tokens in the sequence; consequentially, scaling the compute associated with the inference cannot directly reduce this overhead when using standard decoding schemes. In recent literature studying how to accelerate large language model generation, two primary approaches have been explored: (1) advanced decoding implementations that better parallelize token generation and (2) non-autoregressive language models allowing full sequences to be generated simultaneously.
@@
Speculative decoding accelerates autoregressive generation by leveraging a smaller autoregressive models of the same architecture (the drafter model) to predict candidate sequences, which the original model (the target model) then verifies [12, 24]. Notably, the earliest literature on speculative diffusion adapted a non-autoregressive model to act as the drafter model [11], using a masked language model with a bidirectional decoder [25]. However, the integration of non-autoregressive draft models has not received much attention due to the difficulty introduced by the necessary additional training in existing approaches and the modest speedups that were previously reported using these methods (less than 2x speedup over vanilla decoding schemes).
Thus, recent advancements in speculative decoding have focused on overcoming memory-related constraints, with improvements achieved through various approaches: drafting directly with the target model [26, 27], enhancing draft algorithms [28], and introducing additional parallelization techniques that incorporate branching to refine the drafting process [14, 15, 16].
@@
Models which stray from the autoregressive paradigm have been shown to speedup generation by generating blocks or even entire sequences simultaneously. Gloeckle et al. [20] propose a method of adapting traditional autoregressive models to sample blocks of tokens, improving inference time over similarly scaled models. In a similar vein, diffusion language models have been recognized for their efficiency in generating extended token sequences concurrently, offering even greater speed enhancements. These models recast language generation as a diffusion process either across the embedding space [19] or, more recently, through the probability distributions of generated tokens [18, 17]. Recent models report up to a 32x speedup over similarly sized GPT-2 models [17], and current state-of-the-art further improves runtime speed [18]. However, despite having been shown to dramatically accelerate the inference time for language generation, diffusion models typically perform less effectively than state-of-the-art autoregressive models in terms of standard language metrics, often exhibiting significantly higher perplexity scores [29]. In the following section, we will demonstrate, for the first time, how the speed of these models can be leveraged without being subject to this critical limitation.
@@ 3 Preliminaries and Settings
This process follows a draft-then-verify approach [30], where $M_{q}$ efficiently computes a candidate sequence of tokens, which $M_{p}$ then verifies in parallel.
@@
Previous literature quantifies the likelihood of token acceptance, denoted $\alpha$, and theoretically demonstrate that $\alpha=1-\mathbb{E}(D_{LK}(p,q))$ where $D_{LK}$ represents a divergence measure between distributions [12]. This has led to the prevalent use of drafters taken from the same series as the target models, a paradigm that we challenge in this paper.
@@ 4 Speculative Diffusion Models
Leviathan et al. [12] has conducted theoretical analysis on how to best optimize the value of $\gamma$, however, it has been contingent upon accurately estimating the percentage of tokens in a the sequence that will be accepted by the target model. By their own acknowledgment, it would be necessary to predict this value for each draft and numerically solve for the optimal value of $\gamma$ to fully realize the potential speedup of speculative decoding. Thus, a significant portion of the residual suboptimality in current implementations can be attributed directly to the sensitivity of this hyperparameter.
Diffusion language models are juxtaposed to conventional language models in that they do not sample token sequences in a sequential manner, rather generating entire sequences in parallel. This has resulted in significant speedup over similarly sized autoregressive models when generating extended sequences [17]. This can particularly be observed in longer sequence generations as scaling the draft length $\gamma$ results in minimal overhead due to the ability to directly parallelize token generation.
@@ 4.1 SpecDiff: Formulation
Diffusion models generally operate on continuous data spaces by progressively adding Gaussian noise to the data in a forward diffusion process and then learning to reverse this process to generate new samples [31, 32, 33]. This framework is well-suited for data types such as images, where pixel values can be treated as continuous and thus can naturally accommodate additive Gaussian noise. However, when dealing with discrete data like natural language (tokens), the assumption of continuous noise addition does not hold, as it would result in non-integer values that do not correspond to valid tokens.
Discrete diffusion models address this limitation by redefining the forward processes to transition between discrete token states, such as replacing tokens according to a transition matrix or introducing randomness through categorical distributions [19, 34, 35]. This process enables the generation of coherent and meaningful sequences in natural language processing tasks. However, unlike continuous diffusion models, traditional score-matching techniques cannot be directly applied to learn discrete diffusion models. Instead, various surrogate losses have been proposed for training.
@@
$L_{\infty}^{\text{NELBO}}=\mathbb{E}_{\tilde{q}}\left[\int_{t=0}^{t=1}\frac{(1-\beta_{t})^{\prime}}{\beta_{t}}\sum_{i}\log\langle x_{\theta}(z^{t}_{i}),x_{i}\rangle\,dt\right]$
where $x_{\theta}(z^{t}_{i})$ represents the model’s estimate of the original token $x_{i}$ at time $t$ given the current noisy state $z^{t}_{i}$, and $\beta_{t}$ denotes the noise schedule controlling the diffusion process. Here, $\tilde{q}$ is the forward noising process of in the masked diffusion, defining the distribution over the noisy latent variable $z_{i}^{t}$, and can be related to the NELBO as described in Equation (8) of 18. The expectation $\mathbb{E}_{\tilde{q}}$ signifies averaging over the possible outcomes of $\tilde{q}$ allowing the model to accound for all possible variations of $z_{t}$. The Noise Schedule Derivative term, $\frac{(1-\beta_{t})^{\prime}}{\beta_{t}}$, represents the rate of change in the noise schedule $\beta_{t}$ over time, and $t$ is a continuous timestep between 0 and 1.
@@
       $q_{i+1,\ldots,i+\gamma}^{t-1}(x)\leftarrow M_{q}([x_{0},\ldots,x_{i}]+[q_{i+1,\ldots,i+\gamma}^{t}(x)],t)$
@@
$p_{i}(x),\ldots,p_{i+\gamma+1}(x)\leftarrow$   $M_{p}(x_{0},\ldots,x_{i}),\ldots,M_{p}(x_{0},\ldots,x_{i+\gamma})$
@@
This loss is directly used for pretraining our draft model. As this process learns to denoise over the probability mass vectors, the output of the draft model is a matrix of $\mathbb{R}^{n\times m}$ where $n=\gamma$ and $m$ is the size of the vocabulary. The candidate sequence can then be generated using standard decoding methods over these probability mass vectors, the logits of which are stored to determine whether to accept each draft token.
@@
5 Evaluation
To empirically evaluate the improvements provided by using SpecDiff, the paper provides an empirical analysis on text summarization and more general text generation, leveraging benchmarks that are common to existing literature. Additionally, we assess the performance of our method against the current state-of-the-art using Spec-Bench, a unified evaluation platform for speculative decoding techniques [36]. All evaluation is conducted on two NVIDIA A100 series GPUs (80GB) using CUDA 12.2. Additionally, FlashAttention [37] is used to optimize the performance in all experiments.
5.1 Comparative Analysis of Speculative Decoders
According to the Work-Depth parallel computation model [38], an algorithm can be represented by a DAG, in which each node is an operation and each edge is a dependency. The longest shortest path in this computational DAG (i.e., the longest dependency chain) is called the depth. It follows that the average parallelism of a computation is the total number of nodes divided by the depth, and that higher depth means fewer opportunities for parallelism. We represent the operations by the FLOPs of the model.
@@ 5.1 Comparative Analysis of Speculative Decoders
SpS [12]1 $C_{p}(\gamma)+C_{q}(\gamma)$ $F_{p}(\gamma)+F_{q}(\gamma)$ $D_{p}+\gamma D_{q}$ $1+\gamma$ $M_{q}$
PLD [39] $C_{p}(\gamma)$ $F_{p}(\gamma)+\mathcal{O}(Pn)$ $D_{p}+\mathcal{O}(n)$ $1+\gamma$ None
Lookahead [14]1 $C_{p}\left(2\gamma(n-1)\right)$ $F_{p}\left(2\gamma(n-1)\right)+F_{p}(1)$ $D_{p}$ + $\mathcal{O}(1)$ $n-1$ None
Medusa-1 [26] $C_{p}(\mathcal{T}_{n})+KC_{q}(\gamma)$ $F_{p}(\mathcal{T}_{n})+KF_{q}(\gamma)$ $D_{p}+D_{q}$ $1+\mathcal{T}_{d}$ $M_{q}$
Medusa-2 [26]2 $M_{p}$
Hydra [40]3 $D_{p}+KD_{q}$ $M_{q}$
EAGLE [41] $C_{p}(25)+C_{q}(25)$ $F_{p}(25)+5F_{q}(5)$ $D_{p}+5D_{q}$ $6$ $M_{q}$
EAGLE-2 [42]4 $C_{p}(\mathcal{T}_{n})+C_{q}(\mathcal{T}_{n})$ $F_{p}(\mathcal{T}_{n})+\mathcal{T}_{d}F_{q}(\frac{\mathcal{T}_{n}}{\mathcal{T}_{d}})$ $D_{p}+\mathcal{T}_{d}D_{q}$ $1+\mathcal{T}_{d}$ $M_{q}$
@@
In Medusa, the prediction tree is set by a number of fixed hyperparameters [26], which we aggregate by using the number of nodes and the depth of the tree $\mathcal{T}_{n}$ and $\mathcal{T}_{d}$. Medusa-2 exhibits the same parallel properties as Medusa-1, but requires a joint fine-tuning of $M_{p}$ along with $M_{q}$. Hydra is an adaptation of Medusa, in which the heads are applied in sequence. It thus only affects the depth of the computations.
EAGLE uses a drafter model that contains the target embedding layer, one autoregressive layer, and the target LM head. The fixed tree size is 5 levels deep and contains 25 nodes [41]. EAGLE-2 [42] dispenses with the fixed tree defined in EAGLE, and the induced dynamic tree used in generation is represented with $\mathcal{T}$.

### v4

@@ Abstract
Speculative Diffusion Decoding:
Accelerating Language Generation through Diffusion
Jacob K. Christopher University of Virginia csk4sr@virginia.edu                        Brian R. Bartoldson Lawrence Livermore National Laboratory bartoldson1@llnl.gov
Tal Ben-Nun Lawrence Livermore National Laboratory talbn@llnl.gov                        Michael Cardei University of Virginia ntr2rm@virginia.edu
Bhavya Kailkhura Lawrence Livermore National Laboratory kailkhura1@llnl.gov                        Ferdinando Fioretto University of Virginia fioretto@virginia.edu
@@ 1 Introduction
As autoregressive language modeling with transformers (Vaswani et al., 2017) is scaled to larger compute levels, performance improves and new capabilities emerge (Kaplan et al., 2020; Brown et al., 2020). Indeed, scaling has been shown to to improve the performance of large language models (LLMs) for a diverse array of tasks, including code generation, question answering, summarization, and many other use cases (Achiam et al., 2023; Gemini Team, 2023; Llama Team, 2024). For instance, models such as LLaMA 3.2 90B (Llama Team, 2024), ChatGPT (OpenAI et al., 2024), Cohere 52B (Ruis et al., 2023), Google’s Gemini-ULTRA (Team et al., 2024) exemplify the ongoing trend of deploying and releasing increasingly large models, enabling broader access and application across various domains.
However, while these desired capability arise, running LLMs in inference mode for millions of users produces burdensome electricity, time, and monetary demands. Many methods exist to mitigate these costs – including sparsity, quantization, and distillation – but they often introduce new tradeoffs, e.g., their application can degrade the performance of the model (Hong et al., 2024).
@@
Notably, since both the drafting model’s speed and its alignment relative to the target model are critical to the success of speculative decoding, simultaneous improvements in each of these areas are necessary to ensure speculative decoding’s relevance to future, more capable target models. For instance, a small GPT-2 (Radford et al., 2019) drafting model could produce drafts that are often rejected by GPT-4 (Achiam et al., 2023), and simply scaling the drafting model to address its weaker generations risks diminishing the speed advantage necessary to speculative decoding’s success. To address this challenge, previous efforts have focused on introducing additional parallelization techniques that incorporate prediction trees and branching to refine the drafting process (Fu et al., 2024; Miao et al., 2023b; Svirschevski et al., 2024). However, the gain in generation efficiency are at the expense of much increased number of operations and/or memory for each generation.
This paper proposes a fundamentally different approach to improve speculative decoding: It proposes to replace the auto-regressive drafter with recently introduced discrete diffusion models (Lou et al., 2024; Sahoo et al., 2024). These models offer several key advantages when used as drafters: Firstly, they provide a smooth trade off between the compute cost of generation and the quality of generation (via the number of reverse diffusion steps). Second, while they have historically struggled relative to traditional language models, recent diffusion models have been shown to require 32×32$\times$32 × fewer function evaluations than autoregressive models to produce text with comparable perplexity (Lou et al., 2024), with more recent works reporting even further speedups (Sahoo et al., 2024). This is a trend which is anticipated to continue. Finally, future advances in diffusion model generation quality are highly aligned with their ability to perform strongly as speculative drafters: as drafted tokens are accepted by the target model at a higher rate, a larger number of proposed drafted tokens becomes optimal from an efficiency/speed point of view, and (unlike sequential drafters) diffusion models can easily accommodate generation of many more tokens since they are able to generate entire sequences in a single step.
@@
    The proposed method ensures that all generations from the diffusion language model, which are empirically shown to produce outputs with significantly higher perplexity than current state-of-the-art autoregressive models (Sahoo et al., 2024; Lou et al., 2024; Austin et al., 2021; Gloeckle et al., 2024), align with the outputs generated by larger, more computationally demanding models.
@@
    Finally, the paper sets a new benchmark for speed in language completion tasks on the CNN/DM Nallapati et al. (2016)and OpenWebText datasets Gokaslan et al. (2019).
@@ 2 Related Work
While autoregressive language models provide state-of-the-art performance on language generation tasks, the incremental decoding used by these architectures results in significant overhead at inference time (Miao et al., 2023a). This is largely a result of the inability to parallelize the sequential process of generating tokens in the output sequence as each token generation is dependent upon the preceding tokens in the sequence; consequentially, scaling the compute associated with the inference cannot directly reduce this overhead when using standard decoding schemes. In recent literature studying how to accelerate large language model generation, two primary approaches have been explored: (1) advanced decoding implementations that better parallelize token generation and (2) non-autoregressive language models allowing full sequences to be generated simultaneously.
@@
Speculative decoding accelerates autoregressive generation by leveraging a smaller autoregressive models of the same architecture (the drafter model) to predict candidate sequences, which the original model (the target model) then verifies (Leviathan et al., 2023; Chen et al., 2023). Notably, the earliest literature on speculative diffusion adapted a non-autoregressive model to act as the drafter model (Xia et al., 2023), using a masked language model with a bidirectional decoder (Ghazvininejad et al., 2019). However, the integration of non-autoregressive draft models has not received much attention due to the difficulty introduced by the necessary additional training in existing approaches and the modest speedups that were previously reported using these methods (less than 2x speedup over vanilla decoding schemes).
Thus, recent advancements in speculative decoding have focused on overcoming memory-related constraints, with improvements achieved through various approaches: drafting directly with the target model (Cai et al., 2024; Zhang et al., 2024), enhancing draft algorithms (Sun et al., 2024), and introducing additional parallelization techniques that incorporate branching to refine the drafting process (Fu et al., 2024; Miao et al., 2023b; Svirschevski et al., 2024).
@@
Models which stray from the autoregressive paradigm have been shown to speedup generation by generating blocks or even entire sequences simultaneously. Gloeckle et al. (2024) propose a method of adapting traditional autoregressive models to sample blocks of tokens, improving inference time over similarly scaled models. In a similar vein, diffusion language models have been recognized for their efficiency in generating extended token sequences concurrently, offering even greater speed enhancements. These models recast language generation as a diffusion process either across the embedding space (Austin et al., 2021) or, more recently, through the probability distributions of generated tokens (Sahoo et al., 2024; Lou et al., 2024). Recent models report up to a 32x speedup over similarly sized GPT-2 models Lou et al. (2024), and current state-of-the-art further improves runtime speed Sahoo et al. (2024). However, despite they have been shown to dramatically accelerate the inference time for language generation, diffusion models typically perform less effectively than state-of-the-art autoregressive models in terms of standard language metrics, often exhibiting significantly higher perplexity scores Zheng et al. (2024). In the following section, we will demonstrate, for the first time, how the speed of these models can be leveraged without being subject to this critical limitation.
@@ 3 Preliminaries and Settings
This process follows a draft-then-verify approach (Stern et al., 2018), where $M_{q}$ efficiently computes a candidate sequence of tokens, which $M_{p}$ then verifies in parallel.
@@
Previous literature quantifies the likelihood of token acceptance, denoted $\alpha$, and theoretically demonstrate that $\alpha=1-\mathbb{E}(D_{LK}(p,q))$ where $D_{LK}$ represents a divergence measure between distributions Leviathan et al. (2023). This has led to the prevalent use of drafters taken from the same series as the target models, a paradigm that we challenge in this paper.
@@ 4 Speculative Diffusion Models
Leviathan et al. (2023) has conducted theoretical analysis on how to best optimize the value of $\gamma$, however, it has been contingent upon accurately estimating the percentage of tokens in a the sequence that will be accepted by the target model. By their own acknowledgment, it would be necessary to predict this value for each draft and numerically solve for the optimal value of $\gamma$ to fully realize the potential speedup of speculative decoding. Thus, a significant portion of the residual suboptimality in current implementations can be attributed directly to the sensitivity of this hyperparameter.
Diffusion language models are juxtaposed to conventional language models in that they do not sample token sequences in a sequential manner, rather generating entire sequences in parallel. This has resulted in significant speedup over similarly sized autoregressive models when generating extended sequences (Lou et al., 2024). This can particularly be observed in longer sequence generations as scaling the draft length $\gamma$ results in minimal overhead due to the ability to directly parallelize token generation.
@@ 4.1 SpecDiff: Formulation
Diffusion models generally operate on continuous data spaces by progressively adding Gaussian noise to the data in a forward diffusion process and then learning to reverse this process to generate new samples Sohl-Dickstein et al. (2015b); Ho et al. (2020); Song and Ermon (2019). This framework is well-suited for data types such as images, where pixel values can be treated as continuous and thus can naturally accommodate additive Gaussian noise. However, when dealing with discrete data like natural language (tokens), the assumption of continuous noise addition does not hold, as it would result in non-integer values that do not correspond to valid tokens.
Discrete diffusion models address this limitation by redefining the forward processes to transition between discrete token states, such as replacing tokens according to a transition matrix or introducing randomness through categorical distributions Austin et al. (2021); Hoogeboom et al. (2021); Sohl-Dickstein et al. (2015a). This process enables the generation of coherent and meaningful sequences in natural language processing tasks. However, unlike continuous diffusion models, traditional score-matching techniques cannot be directly applied to learn discrete diffusion models. Instead, various surrogate losses have been proposed for training.
@@
$L_{\infty}^{\text{NELBO}}=\mathbb{E}_{\tilde{q}}\left[\int_{t=0}^{t=1}\frac{(1-\beta_{t})^{\prime}}{\beta_{t}}\sum_{i}\log\langle x_{\theta}(z^{t}_{i}),x_{i}\rangle\,dt\right]$ (1)
where $x_{\theta}(z^{t}_{i})$ represents the model’s estimate of the original token $x_{i}$ at time $t$ given the current noisy state $z^{t}_{i}$, and $\beta_{t}$ denotes the noise schedule controlling the diffusion process. Here, $\tilde{q}$ is the forward noising process of in the masked diffusion, defining the distribution over the noisy latent variable $z_{i}^{t}$, and can be related to the NELBO as described in Equation (8) of Sahoo et al., 2024. The expectation $\mathbb{E}_{\tilde{q}}$ signifies averaging over the possible outcomes of $\tilde{q}$ allowing the model to accound for all possible variations of $z_{t}$. The Noise Schedule Derivative term, $\frac{(1-\beta_{t})^{\prime}}{\beta_{t}}$, represents the rate of change in the noise schedule $\beta_{t}$ over time, and $t$ is a continuous timestep between 0 and 1.
This loss is directly used for pretraining our draft model. As this process learns to denoise over the probability mass vectors, the output of the draft model is a matrix of $\mathbb{R}^{n\times m}$ where $n=\gamma$ and $m$ is the size of the vocabulary. The candidate sequence can then be generated using standard decoding methods over these probability mass vectors, the logits of which are stored to be used when determining whether to accept each draft token.
@@
       $q_{i+1,\ldots,i+\gamma}^{t-1}(x)\leftarrow M_{q}([x_{0:i}]\circ[q_{i+1,\ldots,i+\gamma}^{t}(x)],t)\!\!\!\!$
@@
$p_{i}(x),\ldots,p_{i+\gamma+1}(x)\leftarrow$ $M_{p}(x_{0:i}),\ldots,M_{p}(x_{0:i+\gamma})$
@@
SpS Leviathan et al. (2023)1 $C_{p}(\gamma)+C_{q}(\gamma)$ $F_{p}(\gamma)+F_{q}(\gamma)$ $D_{p}+\gamma D_{q}$ $1+\gamma$ $M_{q}$
PLD Saxena (2023) $C_{p}(\gamma)$ $F_{p}(\gamma)+\mathcal{O}(Pn)$ $D_{p}+\mathcal{O}(n)$ $1+\gamma$ None
Lookahead Fu et al. (2024)1 $C_{p}\left(2\gamma(n-1)\right)$ $F_{p}\left(2\gamma(n-1)\right)+F_{p}(1)$ $D_{p}$ + $\mathcal{O}(1)$ $n-1$ None
Medusa-1 Cai et al. (2024) $C_{p}(\mathcal{T}_{n})+KC_{q}(\gamma)$ $F_{p}(\mathcal{T}_{n})+KF_{q}(\gamma)$ $D_{p}+D_{q}$ $1+\mathcal{T}_{d}$ $M_{q}$
Medusa-2 Cai et al. (2024)2 $M_{p}$
Hydra Ankner et al. (2024)3 $D_{p}+KD_{q}$ $M_{q}$
EAGLE Li et al. (2024b) $C_{p}(25)+C_{q}(25)$ $F_{p}(25)+5F_{q}(5)$ $D_{p}+5D_{q}$ $6$ $M_{q}$
EAGLE-2 Li et al. (2024a)4 $C_{p}(\mathcal{T}_{n})+C_{q}(\mathcal{T}_{n})$ $F_{p}(\mathcal{T}_{n})+\mathcal{T}_{d}F_{q}(\frac{\mathcal{T}_{n}}{\mathcal{T}_{d}})$ $D_{p}+\mathcal{T}_{d}D_{q}$ $1+\mathcal{T}_{d}$ $M_{q}$
@@
5 Comparative Analysis of Speculative Decoders
According to the Work-Depth parallel computation model Blumofe and Leiserson (1999), an algorithm can be represented by a DAG, in which each node is an operation and each edge is a dependency. The longest shortest path in this computational DAG (i.e., the longest dependency chain) is called the depth. It follows that the average parallelism of a computation is the total number of nodes divided by the depth, and that higher depth means fewer opportunities for parallelism. We represent the operations by the FLOPs of the model.
In Medusa, the prediction tree is set by a number of fixed hyperparameters Cai et al. (2024), which we aggregate by using the number of nodes and the depth of the tree $\mathcal{T}_{n}$ and $\mathcal{T}_{d}$. Medusa-2 exhibits the same parallel properties as Medusa-1, but requires a joint fine-tuning of $M_{p}$ along with $M_{q}$. Hydra is an adaptation of Medusa, in which the heads are applied in sequence. It thus only affects the depth of the computations.
EAGLE uses a drafter model that contains the target embedding layer, one autoregressive layer, and the target LM head. The fixed tree size is 5 levels deep and contains 25 nodes Li et al. (2024b). EAGLE-2 Li et al. (2024a) dispenses with the fixed tree defined in EAGLE, and the induced dynamic tree used in generation is represented with $\mathcal{T}$.